See Conceptual Example 6 to review the concepts involved in this problem. A 8.95-kg monkey is hanging by one arm from a branch and swinging on a vertical circle. As an approximation, assume a radial distance of 78.9 cm is between the branch and the point where the monkey's mass is located. As the monkey swings through the lowest point on the circle, it has a speed of 2.72 m/s. Find (a) the magnitude of the centripetal force acting on the monkey and (b) the magnitude of the tension in the monkey's arm.
Answer #1

here,

radial distance , r = 0.789 m

mass of monkey , m = 8.95 kg

speed at the lowest point , v =2.72 m/s

a)

the magnitude of centripital force , F = m * v^2 /r

F = 8.95 * 2.72^2 /0.789

F = 83.9 N

b)

the magnitude of tension in the monkey's arm , T = F + m * g

T = 83.9 + 8.95 * 9.8

T = 171.6 N